OpenStudy (anonymous):

A rectangular storage container with an open top is to have a volume of 10cm^3. The length of its base is twice the width. Material for the base costs 10 dollars per square meter. Material for the sides costs 5 dollars per square meter. Solve for l in terms of w, solve for h in terms of w, and express the cost as a function of w alone.

OpenStudy (campbell_st):

ok... so the dimensions are V = length x width x height and you know length = 2 x width and you know the volume then 10 = 2w x w x h so \[5 = w^2 \times h\] now looking at surface areas.... base = length x width or \[Base = 2w \times w = 2w^2\] front and back A = 2( length x height) \[A = 2( 2w \times h)\] left and right sides A = width x height \[A = w \times h\] so the total surface area is \[SA = 2w^2 + 2wh + wh\] use the volume equation at the top, make h the subject and then subsitute into the surface area equation. so the cost function is Cost = base * 10 + (left and right) *5 + (front + back)*5 or \[Cost = 2w^2 \times 10 + 2(2w \times h) \times 5 + 2(w \times h) \times 5\] hope this helps

OpenStudy (anonymous):

Thanks!